OpenStudy (anonymous):

The zero of F-1(x) in F(x) = x + 3 is A.) -3 B.) 0 C.) 3

mathslover (mathslover):

x+3 = 0 , x=-3 is this what we have to do @hartnn ?

hartnn (hartnn):

zero of F-1(x) means to equate F-1(x) =0 so what is F(0) = ? just put x=0 in F(x)

OpenStudy (anonymous):

Can you please work it out without giving the answer? Im still a bit confused.

hartnn (hartnn):

ok, to get zeros of F-1(x) Equating F-1(x) =0 so, F(0) = 0+3 (put x=0 in F(x)) = ? hence zeros of F-1 (x) = ?

OpenStudy (anonymous):

3

hartnn (hartnn):

yup

OpenStudy (anonymous):

Thanks, I think I get it now!!

hartnn (hartnn):

welcome ^_^ you can also actually calculate F-1 (x) (let it be y) y=x+3 interchange x and y x=y+3 y=x-3 s0 F-1(x) = x-3 to get zeros, put F-1(x)=0 x-3=0 x=3

OpenStudy (anonymous):

Ok thanks...this way seems easier!!!